OpenStudy (anonymous):

how do you get the ? in (3x-7/12-3x) (3x-12/?)=1

OpenStudy (maheshmeghwal9):

\[\frac{3x-7}{12-3x} \times \frac{3x-12}{?}=\frac{3x-7}{12-3x} \times \frac{-1(12-3x)}{?}=\frac{3x-7}{\cancel {12-3x}}\times \frac{-1\cancel{(12-3x)}}{?}=\frac{7x-3}{?}=1.\]

OpenStudy (maheshmeghwal9):

\[\implies \frac{7-3x}{?}=1 \implies ?=\frac{7-3x}{1}=7-3x.\]

OpenStudy (maheshmeghwal9):

Any doubt?

OpenStudy (anonymous):

why did it become 7-3x?

OpenStudy (maheshmeghwal9):

Becoz...... \[(3x-7)\times (-1)= (7-3x).\]

OpenStudy (anonymous):

owwww

OpenStudy (maheshmeghwal9):

now tell me if any doubt.

OpenStudy (anonymous):

thanks--i have no doubt,,thanks again--you are a genius

OpenStudy (maheshmeghwal9):

I omg ---no genius:) yw:)